# Hilbert’s Theorem 90, periodicity, and roots of Artin-Schreier polynomials

### v2

@@ Abstract
Let $E/F$ be a cyclic field extension of degree $n$, and let $\sigma$ generate the group $\textup{Gal}(E/F)$. If $\textup{Tr}^{E}_{F}(y)=\sum_{i=0}^{n-1}\sigma^{i}y=0$, then the additive form of Hilbert’s Theorem 90 asserts that $y=\sigma x-x$ for some $x\in E$. When $E$ has characteristic $p>0$ we prove that $x$ gives rise to a periodic sequence $x_{0},x_{1},\dots$ which has period $pn_{p}$, where $n_{p}$ is the largest $p$-power that divides $n$. We also show, if $y$ lies in the finite field ${\mathbb{F}}_{p^{n}}$, then the roots of a reducible Artin-Schreier polynomial $t^{p}-t-y$ have the form $x+u$ where $u\in{\mathbb{F}}_{p}$ and $x=\sum_{i=0}^{n-1}\sum_{j=0}^{i-1}z^{p^{j}}y^{p^{i}}$ for some $z\in{\mathbb{F}}_{p^{e}}$ with $e=n_{p}$. Furthermore, the sequence $\left(\sum_{j=0}^{i-1}z^{p^{j}}\right)_{i\geqslant 0}$ is periodic with period $pe$. Keywords: Hilbert’s Theorem 90, Artin-Schreier polynomials, periodic sequence 2020 Mathematics Subject Classification: 11T06, 12F10, 11T55
@@ 1 Introduction
Given a quadratic equation $ax^{2}+bx+c=0$ over a finite field ${\mathbb{F}}_{2^{n}}$ of order $2^{n}$, Chen [1] gives ‘formulas’ for the roots that involve $a,b,c$ and some element of trace 1. We give a unified treatment of these formulas which in prime characteristic $p$ depends on $n_{p}$, the largest $p$-power divisor of $n$. This is related the additive version of Hilbert’s Theorem 90 and to sequences of period $pn_{p}$. We say that a sequence $(x_{i})_{i\geqslant 0}$ has period $e$ if $x_{i}=x_{j}$ whenever $i\equiv j\pmod{e}$, and $e$ is minimal. In particular, we prove that a root of a reducible quadratic $x^{2}+x+y\in{\mathbb{F}}_{2^{n}}[x]$ has the form $x=\sum_{i=0}^{n-1}x_{i}y^{2^{i}}$, $x_{i}\in{\mathbb{F}}_{2^{n}}$, where the sequence $(x_{i})_{i\geqslant 0}$ has period $2n_{2}$. If $n_{2}=1$, we may take $x_{i}=0,1$ whenever $i\equiv 0,1\pmod{2}$, and if $n_{2}=2$, we may take $x_{i}=0,\omega,1,\omega^{2}$ whenever $i\equiv 0,1,2,3\pmod{4}$: here $\omega^{2}+\omega+1=0$ and $\omega$ is a primitive cube root of 1. The periodic sequence $(x_{i})_{i\geqslant 0}$ is described in detail, and for larger values of $n_{2}$, in Lemma 4.7.
@@
Let $E/F$ be a cyclic extension of degree $n$. Suppose $\textup{Gal}(E/F)=\langle\sigma\rangle$ and $y,z\in E$ satisfy $\textup{Tr}^{E}_{F}(y)=0$, $\textup{Tr}^{E}_{F}(z)=1$, then $y=\sigma x-x$ where $x=\sum_{i=0}^{n-1}\left(\sum_{j=0}^{i-1}\sigma^{j}z\right)\sigma^{i}y$.
@@
Let $E/F$ be a cyclic extension of degree $n$ where $\textup{Gal}(E/F)=\langle\sigma\rangle$ and $\textup{char}(F)=p$. Let $n$ have $p$-part $n_{p}$, and $p^{\prime}$-part $n_{p^{\prime}}=n/n_{p}$. If $y\in E$ and $\textup{Tr}^{E}_{F}(y)=0$, then
@@
Polynomials of the form $t^{p}-t-y$ where $y\in E$ were studied by Artin and Schreier in the context of generalizing Galois theory to characteristic $p$. Our theorems relate to roots of these Artin-Schreier polynomials. Suppose $\textup{char}(E)=p>0$ and the $p$th power map $\sigma\colon x\mapsto x^{p}$ of $E$ generates $\textup{Gal}(E/F)$, e.g. when $E={\mathbb{F}}_{p^{n}}$ and $F={\mathbb{F}}_{p}$. If $\textup{Tr}^{E}_{F}(y)=0$ and $\textup{Tr}^{E}_{F}(z)=1$, then $R(y,z)=\sum_{i=0}^{n-1}\sum_{j=0}^{i-1}z^{p^{j}}y^{p^{i}}$ is a root of the polynomial $t^{p}-t-y$. For example, if $n_{p}=1$, we may take $z=\frac{1}{n}$ and $x_{i}\coloneq\sum_{j=0}^{i-1}z^{p^{j}}=\frac{i}{n}$ defines a sequence of period $p$, and $x=\sum_{i=0}^{n-1}\frac{i}{n}y^{p^{i}}$ is a root of $t^{p}-t-y$. Because different choices of $n_{p}$ involve different choices for $z$ with $\textup{Tr}^{E}_{F}(z)=1$, this suggests that there are infinitely many (polynomial) formulas for the roots of polynomials $t^{p}-t-y\in E[t]$. If $E={\mathbb{F}}_{p^{n}}$, $F={\mathbb{F}}_{p}$ and $e=n_{p}$, then finding $z$ in ${\mathbb{F}}_{p^{e}}$ with $\textup{Tr}^{E}_{F}(z)=1$ is easy in practise: with probability $1-\frac{1}{p}$ a random element $\zeta\in{\mathbb{F}}_{p^{e}}$ has $\tau\coloneq\textup{Tr}^{E}_{F}(\zeta)\neq 0$, so $z=\tau^{-1}\zeta$ has $\textup{Tr}^{E}_{F}(z)=1$.
There are two versions of Hilbert’s Theorem 90: the additive form and the multiplicative form: we abbreviate these AH90 and MH90, respectively. Theorem 1.1 is different from usual version of AH90 in that the formula for $x$ is very slightly different. This difference leads to Theorem 1.2. We remark first that MH90 is so numbered because it appears as Satz 90 in [2] on p. 272, and Hilbert never actually proved AH90. The standard proof of AH90 is given in Lang [5, p. 215] and it follows Emil Artin’s ideas. In Theorem 1.1, $x$ equals $R(y,z)\coloneq\sum_{i=0}^{n-1}\sum_{j=0}^{i-1}\sigma^{j}z\sigma^{i}y$ where $z\in E$ satisfies $\textup{Tr}^{E}_{F}(z)=1$ and $y=\sigma x-x$ holds. By contrast, the standard proof of AH90 uses a different element, namely $x^{\prime}=R(z,y)$ and $y=x^{\prime}-\sigma x^{\prime}$ holds. The proofs of AH90 I have found, assume $\textup{Tr}^{E}_{F}(z)\neq 0$ (but renormalizing to get $\textup{Tr}^{E}_{F}(z)=1$ is easy) and show that
@@
We remark that the expression $x^{\prime}=R(z,y)$ is very natural in the context of Galois cohomology. Indeed, this form suggests itself via Emmy Noether’s equations, and these equations lead to analogous proofs of AH90 and MH90 as we explain. Emmy Noether generalized the method used to prove AH90 and MH90 by introducing 1-cocyles, to handle a noncyclic (finite) Galois group $\Gamma\coloneq\textup{Gal}(E/F)$, see [3, Theorems 4.29 and 4.30]. Her result [3, Theorem 4.29] says: given a function $y\colon\Gamma\to E\setminus\{0\}$ satisfying $y_{\beta\gamma}=y_{\beta}\beta(y_{\gamma})$ for $\beta,\gamma\in\Gamma$, then for some $z\in E$, $x=\sum_{\beta\in\Gamma}y_{\beta}\beta(z)$ satisfies $y_{\gamma}=x\gamma(x)^{-1}$ for all $\gamma\in\Gamma$. Her result [3, Theorem 4.30] says: given a function $y\colon\Gamma\to E$ satisfying $y_{\beta\gamma}=y_{\beta}+\beta(y_{\gamma})$ for $\beta,\gamma\in\Gamma$, then for some $z\in E$, $x=\textup{Tr}^{E}_{F}(z)^{-1}\sum_{\beta\in\Gamma}y_{\beta}\beta(z)$ satisfies $y_{\gamma}=x-\gamma(x)$ for all $\gamma\in\Gamma$. When $\Gamma=\langle\sigma\rangle$ is cyclic of order $n$, the 1-cocycle conditions imply that $y_{\sigma^{i}}=y(\sigma y)\cdots(\sigma^{i-1}y)$ for $0\leqslant i<n$ where $y_{1}=y$, and similarly in the additive case $y_{\sigma^{i}}=y+(\sigma y)+\cdots+(\sigma^{i-1}y)$ for $0\leqslant i<n$ where $y_{1}=y$. This gives rise to the multiplicative formula $x=\sum_{i=0}^{n-1}\left(\prod_{j=0}^{i-1}\sigma^{j}y\right)\sigma^{i}z$ and when $\textup{Tr}^{E}_{F}(z)=1$, the additive formula $x=R(z,y)\coloneq\sum_{i=0}^{n-1}\left(\sum_{j=0}^{i-1}\sigma^{j}y\right)\sigma^{i}z$, respectively. Theorem 1.1 uses the related expression $R(y,z)=\sum_{i=0}^{n-1}\left(\sum_{j=0}^{i-1}\sigma^{j}z\right)\sigma^{i}y$. Lemma 3.1 in some sense measures the asymmetry between $R(y,z)$ and $R(z,y)$ via the equation $R(y,z)+R(z,y)=-\sum_{i=0}^{n-1}\sigma^{i}(yz)$.
@@ 2 Background
This work fits into a very classical setting, namely generalizing Galois theory to characteristic $p$. The following important result follows from the AH90, see [3, Theorem 4.33] and [5, VIII, Theorem 11] for proofs of Theorem 2.1.
@@
The trace map $\textup{Tr}_{F}^{E}(x)=\sum_{\gamma\in\textup{Gal}(E/F)}\gamma(x)$ of a finite Galois extension $E/F$, is an $F$-linear map $E\to F$ which is nonzero by Artin’s Lemma [5, VIII, Theorem 7]. In certain cases, the trace map can be used to distinguish whether or not $t^{p}-t+y$ is irreducible.
@@
Let $E/F$ be a cyclic Galois extension where $|E:F|=\textup{char}(F)=p$. Suppose that $\textup{Gal}(E/F)=\langle\sigma\rangle$ where $\sigma\colon E\to E$ is the $p$th power map $\sigma(x)=x^{p}$. For $y\in E$, the polynomial $t^{p}-t+y$ is reducible over $E$ if and only if $\textup{Tr}_{F}^{E}(y)=\sum_{i=0}^{p-1}y^{p^{i}}=0$. Furthermore, $E=F(z)$ for some $z\in E$ where $z^{p}-z^{p-1}+1=0$ and $\textup{Tr}^{E}_{F}(z)=1$.
@@
If $\textup{Tr}_{F}^{E}(y)=0$, then $y=z-\sigma(z)=z-z^{p}$ for some $z\in E$ by AH90. Hence $t^{p}-t+y$ is reducible over $E$. Conversely, suppose if $t^{p}-t+y$ is reducible over $E$, then $t^{p}-t+y$ has a root $z\in E$, by Theorem 2.1(b). Using $\sigma z=z^{p}$ shows
$0=\textup{Tr}^{E}_{F}(0)=\textup{Tr}^{E}_{F}(z^{p}-z+y)=\textup{Tr}^{E}_{F}(\sigma z)-\textup{Tr}^{E}_{F}(z)+\textup{Tr}^{E}_{F}(y)=\textup{Tr}^{E}_{F}(y).$
Finally, set $y=1$. Then $\textup{Tr}_{F}^{E}(1)=p\cdot 1=0$ and some $\zeta\in E$ satisfies $t^{p}-t+1=0$. However, $\sigma\zeta=\zeta^{p}=\zeta-1$ implies $\sigma^{i}\zeta=\zeta-i$ so $\zeta$ has $p$ distinct conjugates in $E$ and $E=F(\zeta)$. Therefore $E=F(z)$ where $z=\zeta^{-1}$ is a root of the (reverse) polynomial $t^{p}(t^{-p}-t^{-1}+1)=t^{p}-t^{p-1}+1$. Since the conjugates $z,\sigma z,\dots,\sigma^{p-1}z$ are pairwise distinct, $t^{p}-t^{p-1}+1$ is also irreducible over $F$. The coefficient of $t^{p-1}$ shows that $\textup{Tr}^{E}_{F}(z)=1$. ∎
@@
[…]

### v3

@@ Abstract
Let $E/F$ be a cyclic field extension of degree $n$, and let $\sigma$ generate the group $\textup{Gal}(E/F)$. If $\textup{Tr}^{E}_{F}(y)=\sum_{i=0}^{n-1}\sigma^{i}y=0$, then the additive form of Hilbert’s Theorem 90 asserts that $y=\sigma x-x$ for some $x\in E$. When $E$ has characteristic $p>0$ we prove that $x$ gives rise to a periodic sequence $x_{0},x_{1},\dots$ which has period $pn_{p}$, where $n_{p}$ is the largest $p$-power that divides $n$. We also show, if $y$ lies in the finite field ${\mathbb{F}}_{p^{n}}$, then the roots of a reducible Artin-Schreier polynomial $t^{p}-t-y$ have the form $x+u$ where $u\in{\mathbb{F}}_{p}$ and $x=\sum_{i=0}^{n-1}\sum_{j=0}^{i-1}z^{p^{j}}y^{p^{i}}$ for some $z\in{\mathbb{F}}_{p^{e}}$ with $e=n_{p}$. Furthermore, the sequence $\left(\sum_{j=0}^{i-1}z^{p^{j}}\right)_{i\geqslant 0}$ is periodic with period $pe$.
Keywords: Hilbert’s Theorem 90, Artin-Schreier polynomials, periodic sequences
2020 Mathematics Subject Classification: 11T06, 12F10, 11T55
@@ 1 Introduction
Galois was initially concerned to determine when the roots of polynomial $f(x)$ over a field $F$ of characteristic zero could be expressed as a formula involving the symbols $+,-,\times,/$ and $n$th roots for $n\geqslant 2$. If $F$ has prime characteristic, then the formulas were adapted to included roots of Artin-Schreier polynomials. This note explores different formulas for the roots, replacing roots of Artin-Schreier polynomials with polynomial expressions.
Given a quadratic equation $ax^{2}+bx+c=0$ over a finite field ${\mathbb{F}}_{2^{n}}$ of order $2^{n}$, Chen [1] gives ‘formulas’ for the roots that involve $a,b,c$ and some element of trace 1. We give a unified treatment of these formulas, which in prime characteristic $p$ depends on $n_{p}$, the largest $p$-power divisor of $n$. This is related to the additive version of Hilbert’s Theorem 90 and to sequences of period $pn_{p}$. We say that a sequence $(x_{i})_{i\geqslant 0}$ has period $e$ if $x_{i}=x_{j}$ whenever $i\equiv j\pmod{e}$, and $e$ is minimal. We prove that a root of a reducible polynomial $x^{p}-x-y\in{\mathbb{F}}_{p^{n}}[x]$ has the form $x=\sum_{i=0}^{n-1}x_{i}y^{p^{i}}$, $x_{i}\in{\mathbb{F}}_{p^{n}}$, where the sequence $(x_{i})_{i\geqslant 0}$ has period $pn_{p}$. In the case that $p=2$ and $n$ is odd, we may take $x_{i}=i\in\{0,1\}$, and if $p=2$ and $n$ is twice odd, we may take $x_{i}=0,\omega,1,\omega^{2}$ whenever $i\equiv 0,1,2,3\pmod{4}$: here $\omega$ is a primitive cube root of 1. The periodic sequence $(x_{i})_{i\geqslant 0}$ is described in detail, for larger values of the 2-part $n_{2}$ of $n$, in Lemma 4.9.
@@
Let $E/F$ be a cyclic extension of degree $n$. Suppose $\textup{Gal}(E/F)=\langle\sigma\rangle$ and $y,z\in E$ satisfy $\textup{Tr}^{E}_{F}(y)=0$, $\textup{Tr}^{E}_{F}(z)=1$, then $y=\sigma x-x$ where $x=\sum_{i=0}^{n-1}\kern-1.1pt\left(\sum_{j=0}^{i-1}\sigma^{j}z\right)\kern-1.1pt\sigma^{i}y$.
@@
Let $E/F$ be a cyclic extension of degree $n$ where $\textup{Gal}(E/F)=\langle\sigma\rangle$ and $\textup{char}(F)=p$. Let $n$ have $p$-part $n_{p}$, so $n_{p}\mid n$ and $pn_{p}\nmid n$. If $y\in E$ and $\textup{Tr}^{E}_{F}(y)=0$, then
@@
Polynomials of the form $t^{p}-t-y$ where $y\in E$ were studied by Artin and Schreier in the context of generalizing Galois theory to characteristic $p$: they are analogues of $p$th roots in characteristic 0. Our theorems relate to roots of these Artin-Schreier polynomials. Suppose $\textup{char}(E)=p>0$ and the $p$th power map $\sigma\colon x\mapsto x^{p}$ of $E$ generates $\textup{Gal}(E/F)$, e.g. when $E={\mathbb{F}}_{p^{n}}$ and $F={\mathbb{F}}_{p}$. If $\textup{Tr}^{E}_{F}(y)=0$ and $\textup{Tr}^{E}_{F}(z)=1$, then $R(y,z)=\sum_{i=0}^{n-1}\sum_{j=0}^{i-1}z^{p^{j}}y^{p^{i}}$ is a root of the polynomial $t^{p}-t-y$. For example, if $n_{p}=1$, we may take $z=\frac{1}{n}$ and $x_{i}\coloneq\sum_{j=0}^{i-1}z^{p^{j}}=\frac{i}{n}$ defines a sequence of period $p$, and $x=\sum_{i=0}^{n-1}\frac{i}{n}y^{p^{i}}$ is a root of $t^{p}-t-y$. Because different choices of $n_{p}$ involve different choices for $z$ with $\textup{Tr}^{E}_{F}(z)=1$, this suggests that there are infinitely many (polynomial) formulas for the roots of polynomials $t^{p}-t-y\in E[t]$. If $E={\mathbb{F}}_{p^{n}}$, $F={\mathbb{F}}_{p}$ and $e=n_{p}$, then finding $z$ in ${\mathbb{F}}_{p^{e}}$ with $\textup{Tr}^{E}_{F}(z)=1$ is easy in practise: with probability $1-\frac{1}{p}$ a uniformly random element $\zeta\in{\mathbb{F}}_{p^{e}}$ has $\tau\coloneq\textup{Tr}^{E}_{F}(\zeta)\neq 0$, so $z=\tau^{-1}\zeta$ has $\textup{Tr}^{E}_{F}(z)=1$.
There are two versions of Hilbert’s Theorem 90: the additive form and the multiplicative form: we abbreviate these AH90 and MH90, respectively. Theorem 1.1 is different from usual version of AH90 in that the formula for $x$ is very slightly different. This difference leads to Theorem 1.2. We remark first that MH90 is so numbered because it appears as Satz 90 in [3] on p. 272, and Hilbert never actually proved AH90. The standard proof of AH90 is given in Lang [6, p. 215] and it follows Emil Artin’s ideas. In Theorem 1.1, $x$ equals $R(y,z)\coloneq\sum_{i=0}^{n-1}\sum_{j=0}^{i-1}\sigma^{j}z\sigma^{i}y$ where $z\in E$ satisfies $\textup{Tr}^{E}_{F}(z)=1$ and $y=\sigma x-x$ holds. By contrast, the standard proof of AH90 uses a different element, namely $x^{\prime}=R(z,y)$ (with $y,z$ swapped) and $y=x^{\prime}-\sigma x^{\prime}$ holds. The proofs of AH90 I have found, assume $\textup{Tr}^{E}_{F}(z)\neq 0$ (but scaling to get $\textup{Tr}^{E}_{F}(z)=1$ is easy) and show that
@@
We remark that the expression $x^{\prime}=R(z,y)$ is very natural in the context of Galois cohomology. Indeed, this form suggests itself via Emmy Noether’s equations, and these equations lead to analogous proofs of AH90 and MH90 as we explain. Emmy Noether generalized the method used to prove AH90 and MH90 by introducing 1-cocyles, to handle a noncyclic (finite) Galois group $\Gamma\coloneq\textup{Gal}(E/F)$, see [4, Theorems 4.29 and 4.30]. Her result [4, Theorem 4.29] says: given a function $y\colon\Gamma\to E\setminus\{0\}$ satisfying $y_{\beta\gamma}=y_{\beta}\beta(y_{\gamma})$ for $\beta,\gamma\in\Gamma$, then for some $z\in E$, $x=\sum_{\beta\in\Gamma}y_{\beta}\beta(z)$ satisfies $y_{\gamma}=x\gamma(x)^{-1}$ for all $\gamma\in\Gamma$. Her result [4, Theorem 4.30] says: given a function $y\colon\Gamma\to E$ satisfying $y_{\beta\gamma}=y_{\beta}+\beta(y_{\gamma})$ for $\beta,\gamma\in\Gamma$, then for some $z\in E$, $x=\textup{Tr}^{E}_{F}(z)^{-1}\sum_{\beta\in\Gamma}y_{\beta}\beta(z)$ satisfies $y_{\gamma}=x-\gamma(x)$ for all $\gamma\in\Gamma$. When $\Gamma=\langle\sigma\rangle$ is cyclic of order $n$, the 1-cocycle conditions imply that $y_{\sigma^{i}}=y(\sigma y)\cdots(\sigma^{i-1}y)$ for $0\leqslant i<n$ where $y_{1}=y$, and similarly in the additive case $y_{\sigma^{i}}=y+(\sigma y)+\cdots+(\sigma^{i-1}y)$ for $0\leqslant i<n$ where $y_{1}=y$. This gives rise to the multiplicative formula $x=\sum_{i=0}^{n-1}\left(\prod_{j=0}^{i-1}\sigma^{j}y\right)\sigma^{i}z$ and when $\textup{Tr}^{E}_{F}(z)=1$, the additive formula $x=R(z,y)\coloneq\sum_{i=0}^{n-1}\left(\sum_{j=0}^{i-1}\sigma^{j}y\right)\sigma^{i}z$, respectively. Theorem 1.1 uses the related expression $R(y,z)=\sum_{i=0}^{n-1}\left(\sum_{j=0}^{i-1}\sigma^{j}z\right)\sigma^{i}y$. The asymmetry between $R(y,z)$ and $R(z,y)$ is measured by the equation $R(y,z)+R(z,y)=-\sum_{i=0}^{n-1}\sigma^{i}(yz)$, see Lemma 3.1 below.
@@ 2 Background
This work fits into a very classical setting, namely generalizing Galois theory to characteristic $p$. Theorem 2.1 below follows from the the additive form (AH90) of Hilbert’s Theorem 90, it is proved in [4, Theorem 4.33] and [6, VIII, Theorem 11].
@@
The trace map $\textup{Tr}_{F}^{E}(x)=\sum_{\gamma\in\textup{Gal}(E/F)}\gamma(x)$ of a finite Galois extension $E/F$, is an $F$-linear map $E\to F$ which is nonzero by Artin’s Lemma [6, VIII, Theorem 7]. In certain cases, the trace map can be used to distinguish whether or not $t^{p}-t-y$ is irreducible.
@@
Let $E/F$ be a cyclic Galois extension where $|E:F|=\textup{char}(F)=p$. Suppose that $\textup{Gal}(E/F)=\langle\sigma\rangle$ where $\sigma\colon E\to E$ is the $p$th power map $\sigma(x)=x^{p}$. For $y\in E$, the polynomial $t^{p}-t-y$ is reducible over $E$ if and only if $\textup{Tr}_{F}^{E}(y)=\sum_{i=0}^{p-1}y^{p^{i}}=0$. Furthermore, $E=F(z)$ for some $z\in E$ where $z^{p}-z^{p-1}+1=0$ and $\textup{Tr}^{E}_{F}(z)=1$.
@@
If $\textup{Tr}_{F}^{E}(y)=0$, then $y=\sigma(z)-z=z^{p}-z$ for some $z\in E$ by AH90. Hence $t^{p}-t-y$ is reducible over $E$. Conversely, suppose if $t^{p}-t-y$ is reducible over $E$, then $t^{p}-t-y$ splits over $E$ by Theorem 2.1(b). Let $z\in E$ be a root. Using $\sigma z=z^{p}$ shows
$0=\textup{Tr}^{E}_{F}(0)=\textup{Tr}^{E}_{F}(z^{p}-z-y)=\textup{Tr}^{E}_{F}(\sigma z)-\textup{Tr}^{E}_{F}(z)-\textup{Tr}^{E}_{F}(y)=-\textup{Tr}^{E}_{F}(y).$
Finally, set y=1y=-1. Then $\textup{Tr}_{F}^{E}(-1)=p(-1)=0$ and by the above argument $t^{p}-t+1=0$ is reducible over $E$. Hence some $\zeta\in E$ satisfies $t^{p}-t+1=0$ by Theorem 2.1(b). However, $\sigma\zeta=\zeta^{p}=\zeta-1$ implies $\sigma^{i}\zeta=\zeta-i$ so $\zeta$ has $p$ distinct conjugates in $E$ and $E=F(\zeta)$. Therefore $E=F(z)$ where $z=\zeta^{-1}$ is a root of the (reverse) polynomial $t^{p}(t^{-p}-t^{-1}+1)=t^{p}-t^{p-1}+1$. Since the conjugates $z,\sigma z,\dots,\sigma^{p-1}z$ are pairwise distinct, $t^{p}-t^{p-1}+1$ is irreducible over $F$. The coefficient of $t^{p-1}$ shows that $\textup{Tr}^{E}_{F}(z)=\sum_{i=0}^{p-1}\sigma^{i}z=1$. ∎
@@
A polynomial over a field of characteristic $p$ is solvable by radicals if its roots can be expressed as a finite combination of $\times,\div,\sqrt[r$]{\ \;} where $r$ is a prime different to $p$, and roots of Artin-Schreier polynomials [6, p. 217]. Note that every irreducible polynomial of degree $n$ over a finite field ${\mathbb{F}}_{q}$, is trivially solvable by radicals because each nonzero element of ${\mathbb{F}}_{q^{n}}$ is a primitive $m$th root of unity for some $m\mid(q^{n}-1)$. It makes sense, therefore, to allow non-radical expressions for the roots of polynomials over finite fields. One such example is $R(y,z)=\sum_{i=0}^{n-1}\sum_{j=0}^{i-1}z^{p^{j}}y^{p^{i}}$, which involves the ‘auxiliary’ element $z$.